1) A project manager is using a weighted factor method to decide which of four projects is best, given the company's limited resources. Each project has 1, 2, and 3 criteria. Their importance weights, and scores are as the follows. Project Greenlight has weights of 1, 2, and 3 with scores of 3, 1, and 1.  Project Runway has importance weights of 1, 2, and 3 with scores of 2, 2, and 1. Project X has importance weights of 1, 2, and 3 with scores of 2, 3, and 2. Project Ilenomit has importance weights of 1, 2, and 3 with scores of 2, 2, and 3. Which project should be chosen?

Project

Criteria

Importance Weight

Score

weighted scores

Greenlight

1

1

3

2

2

1

3

3

1

Runway

1

1

2

2

2

2

3

3

1

X

1

1

2

2

2

3

3

3

2

Ilevomit

1

1

2

2

2

2

3

3

3

1.

the least desirable project is 8

2.

the weighted scores for project Runway is 9

3.

the project manager should choose is X

4.

the weighted scores for project X is 15

2) A project requires an initial investment of $200,000 and will generate cash savings of $75,000 each year for the next five years. What is the payback period (round to whole year)?

Year

Cash Flow

Cumulative

0

($200,000)

1

2

3

1.

the payback period is 3

2.

after second year, the remaining balance for the project is ($25,000)
Answer #1

1.

weighted scores:

For Greenlight : 1*3 + 2*1 + 3*1 = 8

For Runway : 1*2 + 2*2 + 3*1 = 9

For X : 1*2 + 2*3 + 3*2 = 14

For Ilevomit: 1*2 + 2*2 + 3*3 = 15

Option a is false, because weighted score is highest of 15 and not 8.

Option b is true as Runway has 9 weighted score

Option c is false as the project manager should choose greenlight( lowest of 8) instead of X(which has 14)

Option d is false as weighted score for X is 14

2.

Payback Period = Cost of Project / Average Annual Cash Inflows = 200000/75000 = 2.67

can be rounded to 3.

So the payback period is 3 will be true.

After 2 years, it will generate only 150000 which is less than 200000 and left with 50000, so it's false

Please leave a like. Thanks a lot.